OpenStudy (anonymous):

What is the graph of one cycle of a sine curve with the given characteristics? y = sin 3q

OpenStudy (anonymous):

is it a graph you are looking for?

OpenStudy (anonymous):

do these by typing \(y=\sin(3x)\) here http://www.wolframalpha.com/

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dsin%283x%29 you see one cycle goes from \(x=0\) to \(x=\frac{2\pi}{3}\)

OpenStudy (anonymous):

Its a graph.

OpenStudy (anonymous):

ok then it is the graph i sent second

OpenStudy (anonymous):

should only go form \(0\) to \(\frac{2\pi}{3}\) for "one cycle"

OpenStudy (anonymous):

Okay. Im gonna show you the graph i think it is.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no should go further

OpenStudy (anonymous):

should go two third of the way to \(\pi\)

OpenStudy (anonymous):

None of them go far.

OpenStudy (anonymous):

send another that may go further

OpenStudy (anonymous):

A

OpenStudy (anonymous):

you found it

OpenStudy (anonymous):

Okay. thank you!

OpenStudy (anonymous):

yw